# Analysis of Multi-Source Language Training in Cross-Lingual Transfer

### v1

@@ Abstract
The successful adaptation of multilingual language models (LMs) to a specific language-task pair critically depends on the availability of data tailored for that condition. While cross-lingual transfer (XLT) methods have contributed to addressing this data scarcity problem, there still exists ongoing debate about the mechanisms behind their effectiveness. In this work, we focus on one of promising assumptions about inner workings of XLT, that it encourages multilingual LMs to place greater emphasis on language-agnostic or task-specific features. We test this hypothesis by examining how the patterns of XLT change with a varying number of source languages involved in the process. Our experimental findings show that the use of multiple source languages in XLT—a technique we term Multi-Source Language Training (MSLT)—leads to increased mingling of embedding spaces for different languages, supporting the claim that XLT benefits from making use of language-independent information. On the other hand, we discover that using an arbitrary combination of source languages does not always guarantee better performance. We suggest simple heuristics for identifying effective language combinations for MSLT and empirically prove its effectiveness.
@@
Seong Hoon Lim, Taejun Yun, Jinhyeon Kim, Jihun Choi, Taeuk Kim Hanyang University, Sony AI {dmammfl,tj1616,kimjinhye0n,kimtaeuk}@hanyang.ac.kr, jihun.a.choi@sony.com
@@ 1 Introduction
There has been recent interest on multilingual language models (LMs), where multiple languages are used in training an LM, due to its capability of accepting multilingual inputs and achieving similar or even better performance compared to monolingual LMs at natural language processing (NLP) tasks. However, despite all these advantages, fine-tuning multilingual LMs using a specific language and an NLP task hinges on acquiring labeled data tailored to a specific purpose, which is not always possible.
@@
While the standard practice in XLT is to employ only one language as source for transfer, there exist attempts that have investigated the use of multiple source languages Singh et al. (2019); Roy et al. (2020); Kew et al. (2023); Chai et al. (2024); Shaham et al. (2024). These studies reveal that employing multiple source languages in XLT—the concept we refer to as Multi-Source Language Training (MSLT)—leads to performance improvement. However, despite evidences that MSLT has clear advantages, several research questions remain unexplored, including: (1) the specific changes that it induces in the internal states of LMs, (2) the conditions under which it achieves its optimal performance, and (3) the breadth of applications where it can be effectively applied, among others.
In this work, we introduce a series of comprehensive analysis exploring the inner workings of MSLT, along with offering guidance for its effective utilization. Specifically, we first attempt to illustrate the utility of MSLT with a couple of intuitive visualizations, to clarify the underlying reasons for its effectiveness. In addition, our experimental results discover various intriguing findings. Regarding the number of languages involved in MSLT, we observe its positive correlation with performance in general, although the trend tends to plateau as the number of languages exceeds a certain threshold.
@@ 4.4 On the Optimal Number of Languages
WikiAnn |
XNLI |
XCOPA |
XWinograd |
XStoryCloze |
MIN |
76.30 (35) | 70.97 (35) | 48.87 (35) | 48.15 (35) | 51.05 (35) | |
@@
Lang2Vec |
Syntax |
85.69 (3) | 73.76 (3) | 51.57 (4) | 55.92 (7) | 57.42 (2) |
Phonology |
84.82 (8) | 73.39 (7) | 49.88 (27) | 58.21 (1) | 55.12 (18) | |
Inventory |
86.07 (2) | 73.77 (2) | 50.50 (11) | 58.20 (3) | 56.45 (10) | |
Family |
84.82 (8) | 73.39 (7) | 49.88 (27) | 58.21 (1) | 55.12 (18) | |
Geometry |
85.58 (4) | 73.24 (9) | 51.72 (2) | 58.12 (6) | 56.62 (9) | |
MAX |
87.05 (1) | 73.86 (1) | 51.76 (1) | 58.21 (1) | 57.98 (1) |
5 Language Set Composition in MSLT
In the previous section, we showed that MSLT exhibits markedly superior performance on XLT compared to SSLT. Assuming the selection of source languages from a total of languages, there exist possible combinations. Given that can be significantly larger than , fine-tuning with every possible combination of source languages to identify the optimal mix is not time- and resource-efficient. Therefore, it is necessary to select an appropriate combination in a more efficient manner. In this section, we first establish the importance of selecting an effective set of source languages and then proceed to examine various criteria for choosing an optimal source language combination.
@@ 5.2 The Harmony of Languages Matters
We compare the performance of MSLT across various source language combinations. According to our experiments, the magnitude of performance improvement varies greatly depending on the combination used, and in some cases, a poor choice may lead to performance deterioration. Figure 5 represents the upper and lower bounds of the performance of MSLT, along with that of SSLT as a baseline. A notable performance gap is observed between the optimal and poor combinations in most experiments. In certain tasks, e.g., XCOPA and XWinograd, MSLT with the least effective combination fails to outperform SSLT. These findings imply that the combination of source languages in MSLT has a substantial impact on performance, highlighting the need to establish criteria for selecting appropriate source languages.
@@ 5.4 Evaluation on Selection Criteria
The findings suggest that selection criteria based on the statistical properties of Language Models (LMs), referenced in Sections §5.3.1 and §5.3.2, do not substantially improve the performance of XLT. In contrast, selecting source languages based on diversity demonstrates a significant performance gain. This implies that in MSLT, prioritizing diversity within the source language combination is more important than relying on the characteristics of corpora used in pretraining LMs.
@@ 6.1 Which Language is Suitable for MSLT?
6.2 Increasing Diversity is Really Helpful?
@@
Experimental results in Figure 6 reveal that constructing combinations with maximum diversity (as in §5.3.3) yields better performance than combinations having minimum diversity (i.e., high similarity) across the majority of tasks.888Full experimental results are listed in Appendix, Table 6.. Notably, there is a significant performance gap between the two contrasting strategies for WikiAnn. We also observe similar patterns for other datasets, confirming the desirability of our decision to consider the diversity of source languages in §5.3.3.
@@ 7 Conclusion
Given the practical availability of language-specific datasets and the ease of experimentation, this paper designates the top 7 languages that appeared most frequently in various multilingual datasets as source language candidates. By analyzing the permutations of selecting three out of these 7 languages, we aim to provide insights based on a wide range of scenarios to establish relevance in a generalized setting. However, it’s important to note that further experiments with a greater number of languages as candidates for source languages are left as future work.
@@
In this study, we utilize models and datasets publicly available on Huggingface, ensuring that no ethical issues are associated with their usage. Additionally, we endeavor to minimize content related to social biases. However, we acknowledge the possibility that some biases may have been ingrained during the pretraining process of language models, highlighting the need for sufficient consideration in this regard.
@@ Appendix A Dataset
XCOPA, known as Cross-lingual Choice of Plausible Alternatives, is a benchmark designed to evaluate machine learning models’ capability to transfer commonsense reasoning across different languages. This dataset comprises translations and reannotations of the English COPA, encompassing 11 languages (Estonian (et), Haitian (ht), Indonesian (id), Italian (it), Quechua (qu), Swahili (sw), Tamil (ta), Thai (th), Turkish (tr)) from diverse language families and regions worldwide. All languages in XCOPA are evaluated as target languages by instruction-tuned models.
@@
XStoryCloze comprises the English StoryCloze dataset professionally translated into 10 non-English languages. Out of these, we choose six languages as target languages, ensuring they did not overlap with the source language set. The selected target languages are as follows: Basque(eu), Hindi (hi), Indonesian (id), Burmese (my), Swahili (sw), and Telugu (te).
@@ Appendix B Training Details
We report the hyperparameters used for training each model in Table 5. For XLM-R and XLM-R, we conducted a hyperparameter search to find an appropriate learning rate for training. As for BLOOM-7B, we attempted to follow the settings from the original lora-tuned model999https://huggingface.co/MBZUAI/bactrian-x-bloom-7b1-lora, but due to the lack of computation resources, we were unable to replicate the original settings such as batch size 128. Instead, we adjusted the training steps the same as the original lora-tuning setting. For BLOOM model, prompt strategies employed for instruciton-tuning and inference are listed in Table 7 and Table 8, respectively.
@@
Common | cutoff_length : 512 & weight decay 0.01 | ||
@@
Method | WikiAnn | XNLI | XCOPA | XWinograd | XStoryCloze | |||||
Most | Least | Most | Least | Most | Least | Most | Least | Most | Least | |
Embedding | 77.93 | 85.58 | 72.88 | 72.26 | 48.85 | 48.85 | 52.61 | 53.33 | 55.12 | 56.98 |
L2V-Syn | 78.52 | 85.69 | 72.87 | 73.76 | 50.32 | 51.57 | 58.21 | 55.92 | 54.05 | 57.42 |
L2V-Pho | 76.68 | 84.82 | 72.88 | 73.39 | 49.88 | 49.88 | 58.20 | 58.21 | 51.05 | 55.12 |
L2V-Inv | 78.71 | 86.07 | 73.13 | 73.77 | 50.85 | 50.50 | 52.98 | 58.20 | 56.92 | 56.45 |
L2V-Fam | 82.95 | 84.82 | 71.90 | 73.39 | 50.75 | 49.88 | 58.20 | 58.21 | 53.62 | 55.12 |
L2V-Geo | 78.71 | 85.58 | 73.13 | 73.24 | 50.85 | 51.72 | 52.98 | 58.12 | 56.92 | 56.62 |
@@
Only Instruction: |
### Input: |
What is the capital of France? |
@@
The capital city of France is Paris. |
Instruction with input: |
### Input: |
Evaluate this sentence for spelling and grammar mistakes. |
He finished his meal and left the restaurant. |
### Output: |
XCOPA |
### Input: |
Answer the number of option which is more plausible for the effect of the situation "Mees keeras kraani lahti." |
1. Tualett täitus veega. |
2. Tilast voolas vett. |
### Output: |
XWinograd |
### Input: |
Answer the number of option which is more plausible for the blank. |
"チンパンジーはリナックスを使えなかった。_が動物だからだ。" |
1. チンパンジー |
2. リナックス |
### Output: |
XStoryCloze |
### Input: |
Answer the number of option which is more natural for the end of story. "Rick tumbuh di keluarga bermasalah. Dia tidak pernah menerima dukungan dari keluarga, dan menjadi anggota geng. Tak butuh waktu lama sampai Rick tertembak dalam sebuah perampokan. Peristiwa itu membuatnya insaf." |
1. Kini dia bahagia. |
2. Dia menjadi anggota geng. |
### Output: |

### v2

@@ Abstract
The successful adaptation of multilingual language models (LMs) to a specific language-task pair critically depends on the availability of data tailored for that condition. While cross-lingual transfer (XLT) methods have contributed to addressing this data scarcity problem, there still exists ongoing debate about the mechanisms behind their effectiveness. In this work, we focus on one of the promising assumptions about the inner workings of XLT, that it encourages multilingual LMs to place greater emphasis on language-agnostic or task-specific features. We test this hypothesis by examining how the patterns of XLT change with a varying number of source languages involved in the process. Our experimental findings show that the use of multiple source languages in XLT—a technique we term Multi-Source Language Training (MSLT)—leads to increased mingling of embedding spaces for different languages, supporting the claim that XLT benefits from making use of language-independent information. On the other hand, we discover that using an arbitrary combination of source languages does not always guarantee better performance. We suggest simple heuristics for identifying effective language combinations for MSLT and empirically prove its effectiveness.
@@
Seong Hoon Lim†, Taejun Yun†, Jinhyeon Kim†, Jihun Choi‡, Taeuk Kim∗† †Hanyang University, ‡Sony AI {dmammfl,tj1616,kimjinhye0n,kimtaeuk}@hanyang.ac.kr, jihun.a.choi@sony.com
@@ 1 Introduction
There has been recent interest in multilingual language models (LMs), where multiple languages are used in training an LM, due to its capability of accepting multilingual inputs and achieving similar or even better performance compared to monolingual LMs at natural language processing (NLP) tasks. However, despite all these advantages, fine-tuning multilingual LMs using a specific language and an NLP task hinges on acquiring labeled data tailored to a specific purpose, which is not always possible.
@@
While the standard practice in XLT is to employ only one language as a source for transfer, there exist attempts that have investigated the use of multiple source languages Singh et al. (2019); Roy et al. (2020); Kew et al. (2023); Chai et al. (2024); Shaham et al. (2024). These studies reveal that employing multiple source languages in XLT—the concept we refer to as Multi-Source Language Training (MSLT)—leads to performance improvement. However, despite it is evident that MSLT has clear advantages, several research questions remain unexplored, including: (1) the specific changes that it induces in the internal states of LMs, (2) the conditions under which it achieves its optimal performance, and (3) the breadth of applications where it can be effectively applied, among others.
In this work, we introduce a series of comprehensive analyses exploring the inner workings of MSLT, along with offering guidance for its effective utilization. Specifically, we first attempt to illustrate the utility of MSLT with a couple of intuitive visualizations, to clarify the underlying reasons for its effectiveness. In addition, our experimental results discover various intriguing findings. Regarding the number of languages involved in MSLT, we observe its positive correlation with performance in general, although the trend tends to plateau as the number of languages exceeds a certain threshold.
@@ 4.4 On the Optimal Number of Languages
5 Language Set Composition in MSLT
In the previous section, we demonstrated that MSLT exhibits markedly superior performance on XLT compared to SSLT. Assuming the selection of source languages from a total of languages, there exist possible combinations. Given that can be significantly larger than , fine-tuning with every possible combination of source languages to identify the optimal mix is not time- and resource-efficient. Therefore, it is necessary to select an appropriate combination more efficiently. In this section, we first establish the importance of selecting an effective set of source languages and then proceed to examine various criteria for choosing an optimal source language combination.
@@ 5 Language Set Composition in MSLT
WikiAnn | XNLI | XCOPA | XWinograd | XStoryCloze | ||
MIN | 76.30 (35) | 70.97 (35) | 48.87 (35) | 48.15 (35) | 51.05 (35) | |
@@
Lang2Vec | Syntax | 85.69 (3) | 73.76 (3) | 51.57 (4) | 55.92 (7) | 57.42 (2) |
Phonology | 84.82 (8) | 73.39 (7) | 49.88 (27) | 58.21 (1) | 55.12 (18) | |
Inventory | 86.07 (2) | 73.77 (2) | 50.50 (11) | 58.20 (3) | 56.45 (10) | |
Family | 84.82 (8) | 73.39 (7) | 49.88 (27) | 58.21 (1) | 55.12 (18) | |
Geometry | 85.58 (4) | 73.24 (9) | 51.72 (2) | 58.12 (6) | 56.62 (9) | |
MAX | 87.05 (1) | 73.86 (1) | 51.76 (1) | 58.21 (1) | 57.98 (1) |
@@ 5.2 The Harmony of Languages Matters
We compare the performance of MSLT across various source language combinations. According to our experiments, the magnitude of performance improvement varies greatly depending on the combination used, and in some cases, a poor choice may lead to performance deterioration. Figure 5 represents the upper and lower bounds of the performance of MSLT, along with that of SSLT as a baseline. A notable performance gap is observed between the optimal and worst combinations in most experiments. In certain tasks, e.g., XCOPA and XWinograd, MSLT with the least effective combination fails to outperform SSLT. These findings imply that the combination of source languages in MSLT has a substantial impact on performance, highlighting the need to establish criteria for selecting appropriate source languages.
@@ 5.4 Evaluation on Selection Criteria
The findings suggest that selection criteria based on the statistical properties of language models, referenced in Sections §5.3.1 and §5.3.2, do not substantially improve the performance of XLT. In contrast, selecting source languages based on linguistic diversity demonstrates a significant performance gain. This implies that in MSLT, prioritizing diversity within the source language combination is more important than relying on the characteristics of corpora used in pretraining LMs.
@@ 6.1 Which Language is Suitable for MSLT?
6.2 Is Increasing Diversity Really Helpful?
@@
Experimental results in Figure 6 reveal that constructing combinations with maximum diversity (as in §5.3.3) yields better performance than combinations having minimum diversity (i.e., high similarity) across the majority of tasks.888Full experimental results are listed in Appendix, Table 6. Notably, there is a significant performance gap between the two contrasting strategies for WikiAnn. We also observe similar patterns for other datasets, confirming the desirability of our decision to consider the diversity of source languages in §5.3.3.
@@ 7 Conclusion
Given the practical availability of language-specific datasets and the ease of experimentation, this paper designates the top 7 languages that appeared most frequently in various multilingual datasets as source language candidates. By analyzing the permutations of selecting three out of these 7 languages, we aim to provide insights based on a wide range of scenarios to establish relevance in a generalized setting. However, it is important to note that further experiments with a greater number of languages as candidates for source languages are left as future work.
@@
In this study, we utilize models and datasets publicly available on Huggingface, ensuring that no ethical issues are associated with their usage. Additionally, we endeavor to minimize content related to social biases. However, we acknowledge the possibility that some biases may have been infused during the pretraining process of language models, highlighting the need for sufficient consideration in this regard.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2020-II201373, Artificial Intelligence Graduate School Program(Hanyang University)). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) under the artificial intelligence semiconductor support program to nurture the best talents (IITP-2024-RS-2023-00253914) grant funded by the Korea government(MSIT). This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(*MSIT) (No.2018R1A5A7059549). *Ministry of Science and ICT.
@@ Appendix A Dataset
XCOPA, known as Cross-lingual Choice of Plausible Alternatives, is a benchmark designed to evaluate machine learning models’ capability to transfer commonsense reasoning across different languages. This dataset comprises translations and re-annotations of the English COPA, encompassing 11 languages (Estonian (et), Haitian (ht), Indonesian (id), Italian (it), Quechua (qu), Swahili (sw), Tamil (ta), Thai (th), Turkish (tr)) from diverse language families and regions worldwide. All languages in XCOPA are evaluated as target languages by instruction-tuned models.
@@
XStoryCloze comprises the English StoryCloze dataset professionally translated into 10 non-English languages. Out of these, we chose six languages as target languages, ensuring they did not overlap with the source language set. The selected target languages are as follows: Basque(eu), Hindi (hi), Indonesian (id), Burmese (my), Swahili (sw), and Telugu (te).
@@ Appendix B Training Details
We report the hyperparameters used for training each model in Table 5. For XLM-R and XLM-R, we conducted a hyperparameter search to find an appropriate learning rate for training. As for BLOOM-7B, we attempted to follow the settings from the original LoRA-tuned model,999https://huggingface.co/MBZUAI/bactrian-x-bloom-7b1-lora but due to the lack of computation resources, we were unable to replicate the original settings such as batch size 128. Instead, we adjusted the training steps the same as the original LoRA-tuning setting. For the BLOOM model, prompt strategies employed for instruction-tuning and inference are listed in Table 7 and Table B, respectively.
@@
Common | cutoff_length: 512 & weight decay: 0.01 | ||
@@
Method | WikiAnn | XNLI | XCOPA | XWinograd | XStoryCloze | |||||||||
Most | Least | Most | Least | Most | Least | Most | Least | Most | Least | |||||
Embedding | 77.93 | 85.58 | 72.88 | 72.26 | 48.85 | 48.85 | 52.61 | 53.33 | 55.12 | 56.98 | ||||
L2V-Syn | 78.52 | 85.69 | 72.87 | 73.76 | 50.32 | 51.57 | 58.21 | 55.92 | 54.05 | 57.42 | ||||
L2V-Pho | 76.68 | 84.82 | 72.88 | 73.39 | 49.88 | 49.88 | 58.20 | 58.21 | 51.05 | 55.12 | ||||
L2V-Inv | 78.71 | 86.07 | 73.13 | 73.77 | 50.85 | 50.50 | 52.98 | 58.20 | 56.92 | 56.45 | ||||
L2V-Fam | 82.95 | 84.82 | 71.90 | 73.39 | 50.75 | 49.88 | 58.20 | 58.21 | 53.62 | 55.12 | ||||
L2V-Geo | 78.71 | 85.58 | 73.13 | 73.24 | 50.85 | 51.72 | 52.98 | 58.12 | 56.92 | 56.62 | ||||
@@
Only Instruction: | ### Input: |
What is the capital of France? | |
### Output: | |
The capital city of France is Paris. | |
Instruction with input: | ### Input: |
Evaluate this sentence for spelling and grammar mistakes. | |
He finished his meal and left the restaurant. | |
### Output: | |
XCOPA | ### Input: |
Answer the number of options which is more plausible for the effect of the situation "Mees keeras kraani lahti." | |
1. Tualett täitus veega. | |
2. Tilast voolas vett. | |
### Output: | |
XWinograd | ### Input: |
Answer the number of option which is more plausible for the blank. | |
{CJK}UTF8min"チンパンジーはリナックスを使えなかった。_が動物だからだ。" | |
1. {CJK}UTF8minチンパンジー | |
2. {CJK}UTF8minリナックス | |
### Output: | |
XStoryCloze | ### Input: |
Answer the number of option which is more natural for the end of story. "Rick tumbuh di keluarga bermasalah. Dia tidak pernah menerima dukungan dari keluarga, dan menjadi anggota geng. Tak butuh waktu lama sampai Rick tertembak dalam sebuah perampokan. Peristiwa itu membuatnya insaf." | |
1. Kini dia bahagia. | |
2. Dia menjadi anggota geng. | |